# The Mirage of Performance Gains: Why Contrastive Decoding Fails to Address Multimodal Hallucination

### v2

@@ 3 Contrastive Decoding for Hallucinations
3.1 Vanilla Decoding
We consider a MLLM parametrized by . The model takes as input a textual query and a visual input , where provides contextual visual information to assist the model in generating a relevant response to the textual query. The response is sampled auto-regressively from the probability distribution conditioned on the query and the visual context . Mathematically, this can be formulated as:
3.2 Visual Contrastive Decoding
@@
where is a hyperparameter controlling the strength of the contrastive adjustment. From the adjusted output distribution , various sampling strategies, such as nucleus sampling and beam search, can be used to generate text.
@@
3.3 Instruction Contrastive Decoding
@@
3.4 Self-Introspective Decoding
@@
Building on VCD and ICD, SID recognizes that directly perturbing the entire original input introduces excessive uncertainty and noise, hindering the induction of the desired hallucination effect. To address this, as shown on the far right of Figure 1, SID adjusts the model architecture by retaining only a small subset of image tokens with low attention scores after the early decoder layers. This adaptive mechanism enhances the generation of vision-and-text association hallucinations during auto-regressive decoding. Subsequently, SID isolates these hallucinations from the original probability distribution, leading to the definition of the contrastive distribution as:
@@
Here, represents the MLLM with structural modifications introduced by SID. Additionally, SID incorporates the adaptive plausibility constraint
4 Misleading Performance Improvement
In this section, we highlight two misleading factors contributing to the performance improvement of contrastive decoding methods on the POPE Benchmark: : Unidirectional output adjustment skews the model towards generating more ”Yes” outputs, leading to a balanced distribution in certain datasets. : The adaptive plausibility constraint degrades sampling decoding strategy into greedy search, resulting in deceptively improved outcomes.
4.1 POPE Benchmark
Polling-based Object Probing Evaluation (POPE) (Li et al., 2023c; Schwenk et al., 2022) is an innovative framework for evaluating object hallucinations in MLLMs. It moves beyond caption-based methods by treating hallucination detection as a binary task, asking straightforward Yes-or-No questions like ”Is there a chair in the image?” POPE’s performance is assessed through four metrics: Accuracy, Precision, Recall, and F1-score, providing a robust evaluation of hallucination behavior in MLLMs.
Notably, the POPE benchmark ensures a balanced label distribution across all dataset subsets, with ”Yes” and ”No” samples each constituting 50% of the total.
4.2 Unidirectional Output Adjustment
In this subsection, we illustrate how contrastive decoding algorithms can deceptively enhance the performance of MLLMs on the POPE Benchmark by applying targeted, unidirectional modifications to the output distribution. We begin by evaluating the performance of various contrastive decoding methods on the MSCOCO (Lin et al., 2014) and GQA (Hudson & Manning, 2019) datasets, analyzing both accuracy and the distribution of the model’s responses.
For this study, we selected LLaVA-v1.5-7B as the foundational MLLM, using a greedy search decoding strategy. The experimental results, summarized in Table 1, show that ICD yielded no performance improvements under this strategy. Consequently, discussions of ICD predictions are temporarily excluded, with the focus shifted to VCD and SID.
| Dataset | MSCOCO-Random | GQA-Adversarial | ||
| Method | Acc (%) | Yes (%) | Acc (%) | Yes (%) |
| Greedy | 87.1 | 39.2 | 80.9 | 54.0 |
| VCD | 88.6 | 46.4 | 78.0 | 63.3 |
| ICD | 86.3 | 38.2 | 80.9 | 52.0 |
| SID | 87.9 | 42.3 | 79.9 | 57.8 |
The results reveal that both VCD and SID significantly biased the model’s output distribution toward ”Yes” across both subsets. On the MSCOCO-Random dataset, where the original output distribution was skewed toward ”No,” VCD and SID corrected this imbalance, resulting in a more balanced distribution and improved accuracy. Conversely, for the GQA-Adversarial subset, where the output distribution was already biased toward ”Yes,” these methods intensified the skew, ultimately reducing prediction accuracy.
We further illustrate how model outputs change after applying contrastive decoding methods, providing a clearer understanding of their performance improvements. As shown in Figure 2, the method primarily alters predictions from ”No” to ”Yes,” significantly outpacing the reverse. On the MSCOCO-Random dataset, where the output distribution is initially skewed toward ”No,” this adjustment converts many false negatives into true positives, thereby improving accuracy. Conversely, on the GQA-Adversarial dataset, which is biased toward ”Yes,” these modifications lead to the misclassification of numerous true negatives as false positives, resulting in a performance decline. For more details on prediction shifts, please refer to Appendix B.
To understand why contrastive decoding methods consistently increase the likelihood of a ”Yes” response, we analyzed the output distribution generated from contrastive samples, as shown in Table 2. In Section 3, we proposed that the primary function of contrastive samples is to induce hallucinations, allowing contrastive decoding to subsequently filter out these hallucinated elements from the output distribution. However, the results in Table 2 reveal that this objective was entirely unmet. Most outputs derived from contrastive samples were incorrect—not due to successfully induced hallucinations, but because the model overwhelmingly favored ”No” responses. This severe bias in the output distribution led to a significant decline in accuracy.
| Dataset | MSCOCO-Random | GQA-Adversarial | ||
| Method | Acc (%) | Yes (%) | Acc (%) | Yes (%) |
| Greedy | 87.1 | 39.2 | 80.9 | 54.0 |
| VCD-C | 76.7 | 28.2 | 71.5 | 41.3 |
| SID-C | 79.0 | 23.6 | 74.2 | 43.1 |
Based on the above discussion, the practical performance of contrastive decoding methods is illustrated in the lower section of Figure 1. The output distribution derived from the contrastive inputs is heavily biased toward ”No.” However, the contrastive decoding method suppresses this content in the original output, thereby unilaterally increasing the model’s likelihood of answering ”Yes.” Whether the model’s performance on the dataset improves depends heavily on whether this increased ”Yes” frequency leads to a more balanced output distribution.
However, as the output distribution of MLLMs tends to be biased toward ”No” in most data subsets, contrastive decoding methods still manage to achieve a strong overall performance on the POPE Benchmark. For more details, please refer to Appendix A.
4.3 Sampling Decoding Degradation
Notably, the POPE Benchmark, which requires models to answer ”Yes” or ”No,” functions as a binary classification task. As a result, greedy search is the most suitable decoding strategy, rendering sampling-based methods unjustifiable. As shown in Table 3, experimental results further confirm that greedy search significantly outperforms sampling. However, many contrastive decoding methods report performance improvements using sampling strategies, necessitating a closer examination of these claims.
@@ 4.3 Sampling Decoding Degradation
We revisit the adaptive plausibility constraint introduced in Section 3.2 and formally defined in Equation 3. This constraint ensures that when the model exhibits high confidence in its outputs corresponding to the original input, the candidate pool is refined to retain only high-probability tokens. By incorporating this mechanism into contrastive decoding methods, it aims to mitigate adverse effects by preventing the generation of implausible tokens, while safeguarding the coherence and quality of the generated content.
Figure 3 provides a detailed explanation of why the adaptive plausibility constraint alone significantly improves performance when using the sampling strategy. For the question ”Is there a cell phone in the image?”, LLaVA-v1.5-7B model generates the correct output distribution: Yes: 8.8% and No: 91.2%. Using a greedy search strategy, the model consistently produces the correct answer, ”No.” However, when employing a sampling strategy, there is an 8.8% chance that the model generates the incorrect answer, ”Yes.”
Among the excluded candidates is the option ”Yes.” Consequently, the sampling strategy reduces to a greedy search, ensuring a 100% probability of correctly answering ”No.”
Consequently, the adaptive plausibility constraint greatly limits the pool of candidate options, transforming the sampling strategy into a predominantly greedy search. As demonstrated in Table 3 and Figure 3, MLLMs exhibit markedly superior performance on the POPE Benchmark under a greedy strategy, underscoring the constraint’s pivotal contribution to performance gains.
This also explains why, in Section 4.2, ICD does not enhance model performance under a greedy search. However, in its original paper, where a sampling strategy was employed, the reported performance gains were much more pronounced.
4.4 Insights
In Section 4.2, we show that when using greedy search as the decoding strategy, contrastive decoding methods modify the model’s predictions in a unidirectional manner, shifting the output distribution toward Yes. As a result, performance improvements primarily depend on whether the model’s original output distribution was biased toward No.
In Section 4.3, we demonstrate that when sampling is used as the decoding strategy, the adaptive plausibility constraint effectively reduces it to greedy search, serving as a key driver of the observed performance gains.
5 Spurious Improvement Methods
In this section, we propose a series of spurious improvement methods based on the two fundamental reasons for performance improvement discussed in Section 4.4. Although these methods are entirely unrelated to hallucination mitigation, they yield experimental results comparable to contrastive decoding techniques. This evidence suggests that while contrastive decoding enhances performance on POPE Benchmark, it does not address hallucinations.
5.1 Forced Distribution Adjustment
For the first misleading factor in performance improvement, which involves modifying model predictions in a single direction to bias the output distribution toward ”YES,” we introduce two pseudo-performance enhancement methods: Prompt-Based Adjustment and Output Layer Modification, as illustrated in Figure 4.
Prompt-Based Adjustment modifies the input side of the model by appending an additional prompt, ”Answer Yes if possible,” after the user’s instruction. This extra input biases the model’s output distribution toward ”Yes.”
| LLaVA-v1.5-7B | LLaVA-v1.5-13B | QwenVL-Chat-7B | ||||||||
| Category | Method | Accuracy | F1-Score | Yes(%) | Accuracy | F1-Score | Yes(%) | Accuracy | F1-Score | Yes(%) |
| Greedy | 87.1 | 85.5 | 39.2 | 86.7 | 85.0 | 38.7 | 85.9 | 84.0 | 37.8 | |
| VCD | 88.6 | 88.2 | 46.4 | 89.2 | 88.5 | 44.4 | 87.7 | 86.4 | 40.6 | |
| SID | 87.9 | 86.9 | 42.4 | 87.2 | 86.2 | 42.5 | 86.5 | 85.3 | 39.9 | |
| PBA | 87.6 | 86.3 | 40.2 | 90.2 | 89.7 | 45.7 | 87.3 | 86.2 | 41.5 | |
| Random | OLM | 89.6 | 89.0 | 44.2 | 90.0 | 89.9 | 48.8 | 88.2 | 87.3 | 43.8 |
| Greedy | 85.8 | 84.3 | 40.4 | 86.0 | 84.3 | 39.4 | 85.6 | 83.6 | 38.2 | |
| VCD | 86.2 | 86.0 | 48.8 | 87.3 | 86.8 | 46.3 | 87.1 | 85.9 | 41.2 | |
| SID | 85.1 | 84.4 | 45.1 | 85.1 | 84.3 | 44.6 | 85.3 | 84.2 | 39.8 | |
| PBA | 86.2 | 85.0 | 41.6 | 88.4 | 88.1 | 47.5 | 86.8 | 85.5 | 42.3 | |
| Popular | OLM | 87.3 | 86.9 | 46.5 | 88.6 | 88.6 | 50.2 | 87.4 | 86.7 | 44.8 |
| Greedy | 83.6 | 82.3 | 42.6 | 84.3 | 82.7 | 41.0 | 84.0 | 82.2 | 39.7 | |
| VCD | 81.9 | 82.5 | 53.1 | 83.8 | 83.8 | 49.7 | 84.5 | 83.4 | 43.7 | |
| SID | 82.3 | 82.0 | 47.9 | 82.9 | 82.3 | 46.9 | 83.2 | 82.1 | 42.5 | |
| PBA | 83.7 | 82.7 | 44.0 | 84.5 | 84.7 | 51.3 | 84.1 | 83.2 | 45.2 | |
| Adversarial | OLM | 83.6 | 83.6 | 50.1 | 83.9 | 84.7 | 54.9 | 84.8 | 84.5 | 48.4 |
Output Layer Modification refers to adjustments made at the output stage of the model. After generating its initial prediction, the model evaluates the probabilities of ”Yes” and ”No.” If their difference is small, i.e.,
the prediction is forcibly set to ”Yes.” Here, controls how close the probabilities must be to trigger this modification. This adjustment significantly increases the likelihood of the model predicting ”Yes.”
Experimental Settings. We selected QwenVL-7B, LLaVA-v1.5-7B, and LLaVA-v1.5-13B as the backbone MLLMs. For decoding, we employed a greedy search strategy. Our experiments were conducted on COCO dataset, where the raw output distribution of MLLMs tends to be biased toward ”No.” As a result, all contrastive decoding methods exhibited notable performance improvements on this dataset.
Results and analysis. The experimental results, presented in Table 4, clearly show that PBA achieves an even greater performance improvement than SID, while OLM surpasses VCD. Prediction accuracy increases as the output distribution approaches balance. However, as highlighted in red, when the distribution shifts beyond balance and biases toward ”Yes,” accuracy begins to decline again, perfectly aligning with the conclusion in Section 4.4.
Although PBA and OLM are not designed for hallucination mitigation, they produce results similar to contrastive decoding methods, strongly suggesting that contrastive decoding does not effectively address hallucinations. For the experimental results on the AOKVQA and GQA datasets, please refer to Appendix C.
| Category | Method | LLaVA-v1.5-7B | LLaVA-v1.5-13B | QwenVL-Chat-7B | ||||||
| Accuracy | F1-Score | Yes(%) | Accuracy | F1-Score | Yes(%) | Accuracy | F1-Score | Yes(%) | ||
| Random | sample | 83.8 0.0 | 83.1 | 45.6 | 84.6 0.0 | 83.9 | 45.9 | 81.5 0.0 | 79.6 | 41.1 |
| VCD | 86.6 2.8 | 86.9 | 52.5 | 86.7 2.1 | 86.7 | 49.5 | 83.8 2.3 | 82.7 | 44.0 | |
| ICD | 85.2 1.4 | 84.5 | 47.0 | 85.8 1.2 | 85.0 | 44.9 | 82.5 1.0 | 81.8 | 42.0 | |
| SID | 84.9 1.1 | 84.8 | 49.1 | 86.0 1.4 | 86.0 | 49.8 | 82.9 1.4 | 82.5 | 43.5 | |
| sample† | 85.4 1.6 | 84.7 | 45.1 | 86.1 1.5 | 85.4 | 45.3 | 83.0 1.5 | 82.0 | 41.8 | |
| Popular | sample | 77.3 0.0 | 77.8 | 52.1 | 80.6 0.0 | 80.6 | 49.9 | 76.8 0.0 | 75.9 | 46.1 |
| VCD | 78.7 1.4 | 80.4 | 59.4 | 82.9 2.3 | 83.3 | 52.4 | 78.2 1.4 | 78.1 | 49.4 | |
| ICD | 78.1 0.8 | 79.3 | 54.0 | 81.5 0.9 | 81.3 | 49.3 | 77.5 0.7 | 76.8 | 47.2 | |
| SID | 78.4 1.1 | 79.1 | 53.7 | 82.5 1.9 | 83.0 | 53.3 | 77.9 1.1 | 77.5 | 48.0 | |
| sample† | 78.6 1.3 | 79.6 | 52.0 | 81.8 1.2 | 81.8 | 49.6 | 78.1 1.3 | 77.0 | 46.8 | |
| Adversarial | sample | 75.1 0.0 | 76.1 | 54.1 | 78.2 0.0 | 78.9 | 53.2 | 76.4 0.0 | 75.3 | 45.5 |
| VCD | 76.4 1.3 | 79.1 | 62.5 | 80.3 2.1 | 81.6 | 57.0 | 78.6 2.2 | 78.5 | 49.2 | |
| ICD | 75.8 0.7 | 76.6 | 54.2 | 79.2 1.0 | 79.8 | 52.8 | 76.8 0.4 | 76.1 | 46.0 | |
| SID | 76.3 1.2 | 77.8 | 57.5 | 78.7 0.5 | 80.1 | 57.5 | 77.2 0.8 | 76.9 | 47.5 | |
| sample† | 76.3 1.2 | 77.3 | 54.2 | 79.5 1.3 | 80.1 | 53.1 | 77.9 1.5 | 76.5 | 46.2 | |
5.2 Standalone Application of the Constraint.
Experimental Settings. We utilize LLaVA-v1.5-7B and LLaVA-v1.5-13B as our foundational MLLMs, employing a sampling decoding strategy. Our experiments are conducted on the GQA dataset, where the original output distribution of MLLMs is relatively balanced. Consequently, the modification introduced by contrastive decoding methods, which shifts the output distribution towards ”Yes,” does not introduce a positive bias. However, since the adaptive plausibility constraint converts the sampling strategy into a greedy search, the model’s performance still improves.
Results and analysis. The experimental results in Table 5 show that when MLLMs adopt sampling as the decoding strategy, applying the adaptive plausibility constraint alone yields an approximate 2.5% performance improvement, effectively validating the conclusion in Section 4.4. Notably, since the adaptive plausibility constraint is entirely unrelated to hallucination mitigation yet achieves performance on par with various contrastive decoding methods, this strongly suggests that contrastive decoding methods do not actually mitigate hallucinations. For the experimental results on the AOKVQA and COCO datasets, please refer to Appendix D.
6 Discussion on Hallucination Mitigation
Based on the insights from Sections 4 and 5, we propose some new criteria for evaluating hallucination mitigation.
Impact of Decoding Strategies. When evaluating on POPE Benchmark, it is essential to account for the substantial influence of different decoding strategies on model performance. Notably, greedy search consistently outperforms sampling-based approaches such as nucleus sampling and beam search. If a hallucination mitigation method involves modifications to the sampling module, careful consideration must be given to whether these changes affect the core properties of the decoding strategy.
Balancing Correction and Preservation. An effective hallucination mitigation method must strike a balance: it should correct incorrect answers while preserving correct ones. However, as shown in Figure 2, some flawed approaches, despite fixing many errors, also introduce unnecessary modifications to originally correct responses. This behavior resembles mere answer editing rather than genuine hallucination mitigation. To enhance evaluation rigor, future studies should explicitly report instances where correct responses are mistakenly altered, providing a clearer measure of a method’s true effectiveness.
7 Conclusion
This study demonstrates that the performance improvements of contrastive decoding on the POPE benchmark largely stem from two misleading factors: (1) a unidirectional shift in the model’s output distribution, which biases it toward generating ”Yes” responses, artificially balancing the distribution in certain datasets, and (2) the adaptive plausibility constraint, which reduces sampling decoding to greedy search. By comparing experimental results from spurious improvement methods and contrastive decoding, we confirm that while contrastive decoding enhances performance, it ultimately fails to mitigate hallucinations.
Impact Statement
Appendix A Skewness in the Output Distribution
As shown in Figure 5, the raw output distribution of MLLMs across different datasets exhibits skewness, with most datasets displaying a tendency for MLLMs to favor ”No” in their outputs.
Appendix B Additional Research on Changes in Prediction
After applying Visual Contrastive Decoding, the prediction shifts of the LLaVA-v1.5-7B model are shown in Figure 6. It is evident that across all datasets, the number of samples transitioning from Positive to Negative is significantly smaller than those shifting from Negative to Positive. This indicates that the model’s output distribution is biased towards Yes.
Appendix C Further Experiments on Forced Distribution Adjustment
Tables 6 and 7 present the performance of Prompt-Based Adjustment (PBA) and Output Layer Modification (OLM) on the AOKVQA and GQA datasets. On both datasets, PBA and OLM maintain performance levels comparable to contrastive decoding methods. However, after surpassing the balanced distribution, accuracy declines rather than improving.
@@ Appendix C Further Experiments on Forced Distribution Adjustment
Appendix D Further Experiments on Standalone Application of the Constraint
Tables 8 and 9 present the performance of the adaptive plausibility constraint on the AOKVQA and COCO datasets. When applied independently, the adaptive plausibility constraint consistently improves performance by 1.5% to 2% across both datasets.

### v3

@@ Appendix A Contrastive Decoding for Hallucinations
A.1 Visual Contrastive Decoding
@@
A.2 Instruction Contrastive Decoding
@@
A.3 Self-Introspective Decoding
@@
Building on VCD and ICD, SID recognizes that directly perturbing the entire original input introduces excessive uncertainty and noise, hindering the induction of the desired hallucination effect. To address this, as shown on the far right of Figure˜1, SID adjusts the model architecture by retaining only a small subset of image tokens with low attention scores after the early decoder layers. This adaptive mechanism enhances the generation of vision-and-text association hallucinations during auto-regressive decoding. Subsequently, SID isolates these hallucinations from the original probability distribution, leading to the definition of the contrastive distribution as:
@@
Here, represents the MLLM with structural modifications introduced by SID. Additionally, SID incorporates the adaptive plausibility constraint.
Appendix B Additional Research on Changes in Prediction
After applying Visual Contrastive Decoding, the prediction shifts of the LLaVA-v1.5-7B model are shown in Figure˜6. It is evident that across all datasets, the number of samples transitioning from Positive to Negative is significantly smaller than those shifting from Negative to Positive. This indicates that the model’s output distribution is biased towards Yes.
Appendix C The True Impact of Contrastive Perturbations on Prediction Outcomes
We begin by presenting our conclusion: image perturbations in visual contrastive decoding can cause the model to randomly "overlook" certain elements of the image due to their stochastic nature. Consequently, the model’s response is based on only a partial view of the input. This phenomenon is illustrated with a concrete example, as shown in Figure˜7.
With the original visual input, the model exhibited a hallucination, responding: “There are four uncut fruits in the image.” After introducing diffusion noise and setting the diffusion time steps to 100, 300, and 500 respectively, the LLaVA-v1.5-7B model produced the following responses: “There are 3 / 2 / 0 uncut fruits in the image,” correspondingly. These results demonstrate that the presence of noise causes the model to lose access to certain localized visual details, leading it to respond based on the remaining visible features.
This also helps explain why, in POPE benchmark, contrastive examples often cause the model to answer “No”, not due to external bias, but because it genuinely fails to detect the relevant object. However, this effect should not be mistaken for hallucination. In fact, contrastive samples do not induce hallucinations. For example, when visual contrastive decoding is applied to this case, the model correctly outputs: “There are four uncut fruits in the image.”
This brings us back to the foundational assumption behind contrastive decoding methods for mitigating hallucinations. These approaches are grounded in the belief that object hallucinations in MLLMs primarily stem from overly dominant linguistic priors, and thus attempt to suppress such priors through contrastive mechanisms. However, in practice, hallucinations often arise because MLLMs fail to correctly ground specific semantic concepts in the visual input, for instance, the concept of “uncut” in the previous example. As such, the failure of visual contrastive decoding can be attributed to a fundamental flaw in its underlying assumption.
Appendix D Performance across different decoding strategies.
Table˜7 presents the impact of various decoding strategies on the prediction performance of LLaVA-v1.5-7B evaluated on the POPE Benchmark. It is evident that the greedy strategy yields significantly better results compared to the sampling strategy.
@@ Appendix C The True Impact of Contrastive Perturbations on Prediction Outcomes
Appendix E Further Experiments on Forced Distribution Adjustment
Tables˜8 and 9 present the performance of Prompt-Based Adjustment (PBA) and Output Layer Modification (OLM) on the AOKVQA and GQA datasets. On both datasets, PBA and OLM maintain performance levels comparable to contrastive decoding methods. However, after surpassing the balanced distribution, accuracy declines rather than improving.
@@ Appendix E Further Experiments on Forced Distribution Adjustment
Appendix F Further Experiments on Standalone Application of the Constraint
Tables˜10 and 11 present the performance of the adaptive plausibility constraint on the AOKVQA and COCO datasets. When applied independently, the adaptive plausibility constraint consistently improves performance by 1.5% to 2% across both datasets.